### v1

@@ 1 Introduction
Our "Sources of capital growth" [^1] Getty and Tkachenko (2024) proposed the term "thrift theory" to mean the commonly accepted doctrine that net saving, within the technological growth rate ("warranted rate"), gives capital growth. "Free growth theory" was defined as an opposite hypothesis that capital growth is explained wholly by a gain in productivity of capital and labor already present, so that net saving is superflous. Data for net saving and market-value capital taken from natural accounts of 92 nations were argued to support free growth theory. Here we pursue a parallel argument and test from data for consumption.
@@ 2 Consumption and net saving
Thrift theory begins with the idea that net saving $S_{net}$, within the warranted rate, gives capital growth $\Delta K$. That is,
$\Delta K=S_{net}\ ,\quad\text{in thrift theory.}$. (1)
@@
$g(K)=s^{*}\quad\text{in thrift theory.}$. (2)
In thrift theory and common teaching, where innovation provides a required condition for capital growth, but only insofar as net income is saved in investment, and where consumption and net saving sum to net income; it follows that any change in net saving implies an equal and opposite change in consumption $C$222i.e., a suspension of consumption provides the only source of a gain in net saving, arguendo, while a gain in consumption becomes the only disposition of net income not applied to net saving. That is,
@@
$\Delta g(K)=\Delta s^{*}=-\Delta c^{*}\ ,\quad\text{from which}\quad\frac{-\Delta c^{*}}{\Delta g(K)}=1\ ,\quad\text{in thrift theory.}$. (4)
@@
Fig. 1: Plots of $\theta_{c}$ averaged over all countries, GDP-weighted. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\ \theta_{c}=1$.
Table 1: Regression of $-\Delta c^{*}$ on $\Delta g(K)$, all countries and years, GDP-weighted. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\text{Regression}=1$.
@@
Table 2: $\theta_{c}$ in Countries and Periods Shown. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\theta_{c}=1.$
@@ 3 Balanced growth
Keynes (1936) taught that households tend to consume a constant proportion of income. Consumption would tend to peak somewhere between a worker’s mid-forties and retirement, then, as income does. Friedman (1957) proposed conversely that households consider expected lifetime income, not current income alone, and meet current consumption needs from accumulated savings as well as current income within that total. Ando and Modigliani (1963) followed this lead by suggesting that younger households tend to save to afford children and retirement, and that older ones dissave to pay consumption costs of both as they arrive.
In any of these scenarios, random distribution of ages could allow phases of more saving in some households to coincide with phases of more consumption in other households. It becomes possible, then, if net saving realizes equal capital growth, for consumption, net saving, capital growth and capital itself all to grow at the same rate. Then net output $Y_{net}$ can grow at that same rate if the capital/output ratio holds constant. This scenario where $C$, $S_{net}$, $K$ and $Y_{net}$ all grow at the same rate is called balanced growth.
@@
Test results are conclusive. $\theta_{c}$ and the regression of $-\Delta c^{*}$ on $\Delta g(K)$ are both predicted at unity in thrift theory. They are measured respectively at near zero and substantially less than zero. These findings parallel and reinforce our similar ones from data for net saving and consumption reported in Getty and Tkachenko (2024).

### v2

@@ 1 Introduction
Our "Sources of capital growth" [^1] Getty and Tkachenko (2024) proposed the term "thrift theory" to mean the commonly accepted doctrine that net saving, within the technological growth rate ("warranted rate"), gives capital growth. "Free growth theory" was defined as an opposite hypothesis that capital growth is explained wholly by a gain in productivity of capital and labor already present, so that net saving is superflous. Data for net saving and market-value capital taken from national accounts of 92 nations were argued to support free growth theory. Here we pursue a parallel argument and test from data for consumption.
@@ 2 Consumption and net saving
Thrift theory begins with the idea that net saving $S_{net}$, within the warranted rate, gives capital growth $\Delta K$. [^2] Getty and Tkachenko (2024) questioned the doctrine that saving equals investment, but equated them in that paper by using the word ”saving” in the Keynesian sense of saving actually invested. We apply the same meaning here. That is,
$\Delta K=S_{net}\ ,\quad\text{in thrift theory.}$ (1)
@@
$g(K)=s^{*}\quad\text{in thrift theory.}$ (2)
In thrift theory and common teaching, where innovation provides a required condition for capital growth, but only insofar as net income is saved in investment, and where consumption and net saving sum to net income; it follows that any change in net saving implies an equal and opposite change in consumption CC.333i.e., a suspension of consumption provides the only source of a gain in net saving, arguendo, while a gain in consumption becomes the only disposition of net income not applied to net saving. That is,
@@
$\Delta g(K)=\Delta s^{*}=-\Delta c^{*}\ ,\quad\text{from which}\quad\frac{-\Delta c^{*}}{\Delta g(K)}=1\ ,\quad\text{in thrift theory.}$ (4)
@@
Fig. 1: Plots of $\theta_{c}$ averaged over all countries, GDP-weighted. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\ \theta_{c}\cong 1$.
Table 1: Regression of $\Delta c^{*}$ on $\Delta g(K)$, all countries and years, GDP-weighted. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\text{Regression}\cong 1$.
@@
Table 2: $\theta_{c}$ in Countries and Periods Shown. Screen: $|\Delta g(K)|\geq 0.01$. $H_{0}:\theta_{c}\cong 1.$
@@ 3 Balanced growth
Keynes (1936) taught that households tend to consume a constant proportion of income. Consumption would tend to peak somewhere between a worker’s mid-forties and retirement, then, as income does. Franco Modigliani, in 1954, [^6] Modigliani and Brumberg (1954) proposed conversely that younger households tend to save to afford children and retirement, and that older ones dissave to pay consumption costs of both as they arrive. Milton Friedman suggested in 1957 that households consider lifetime income, not current income alone, and choose between saving and consumption accordingly.
In any of these three scenarios, random distribution of ages could allow phases of more saving in some households to coincide with phases of more consumption in other households. It becomes possible, then, if net saving realizes equal capital growth, for consumption, net saving, capital growth and capital itself all to grow at the same rate for all households together. Then net output $Y_{net}$ could grow at that same rate if the capital/output ratio holds constant. This scenario where $C$, $S_{net}$, $K$ and $Y_{net}$ all grow at the same rate is called balanced growth.
@@
Test results are conclusive. $\theta_{c}$ and the regression of $\Delta c^{*}$ on $\Delta g(K)$ are both predicted at unity in thrift theory. They are measured respectively at near zero and substantially less than zero. These findings parallel and reinforce our similar ones from data for net saving and consumption reported in Getty and Tkachenko (2024).